A 30 cm3 sample of nitrogen was reacted with a 60 cm3 sample of fluorine according to the equation
What is the volume of the gas mixture after the reaction, at constant temperature and pressure?
1
Expert's answer
2020-07-08T08:24:16-0400

The equation is

N2(g) + 3F2(g) --> 2NF3(g)

The mole (and volume) N2 : F2 ratio according to the equation is 1 : 3.

The actual ratio is 30 : 60 or 1 : 2, hence F2 is the limiting reactant.

To calculate the volume of a gas mixture after the reaction, the volume of NF3 and remaining N2 should be determined.

V(NF3) = 2/3 x V(F2) = 40 cm3.

According to the equation, the volume of N2 that reacted is 1/3 x V(F2) = 20 cm3.

Vremaining(N2) = Vtotal(N2) - Vreacted(N2)= 30 - 20 = 10cm3.

Vmixture = 40 + 10 = 50 cm3.

Answer: 50 cm3